### v1

# Asymmetric REINFORCE for off-Policy Reinforcement Learning:

Abstract
Reinforcement learning (RL) is increasingly used to align large language models (LLMs). Off-policy methods offer greater implementation simplicity and data efficiency than on-policy techniques, but often result in suboptimal performance. In this work, we study the intermediate range of algorithms between off-policy RL and supervised fine-tuning by analyzing a simple off-policy REINFORCE algorithm, where the advantage is defined as $A=r-V$, with $r$ a reward and $V$ some tunable baseline. Intuitively, lowering $V$ emphasizes high-reward samples, while raising it penalizes low-reward ones more heavily. We first provide a theoretical analysis of this off-policy REINFORCE algorithm, showing that when the baseline $V$ lower-bounds the expected reward, the algorithm enjoys a policy improvement guarantee. Our analysis reveals that while on-policy updates can safely leverage both positive and negative signals, off-policy updates benefit from focusing more on positive rewards than on negative ones. We validate our findings experimentally in a controlled stochastic bandit setting and through fine-tuning state-of-the-art LLMs on reasoning tasks.
[…]
Charles Arnal at
1 Introduction
Reinforcement Learning (RL) has long been applied to align Large Language Models (LLMs) to users’ preferences using human feedback (Christiano et al., 2017; Ouyang et al., 2022; Dubey et al., 2024); more recently, it has been used to augment models in more general ways, and in particular to develop their reasoning, coding and tool use capacities (Shao et al., 2024; Guo et al., 2025; Meta, 2025; OpenAI, 2025). As RL theoretically allows models to learn beyond the limits of existing training data (see e.g. Silver et al. (2016); AlphaEvolve-team (2025)), we can expect it to play an increasingly crucial role in years to come as models begin pushing scientific boundaries. So far, on-policy techniques, in which models are trained on samples that they generated, have been preferred to off-policy methods, in which models are trained using samples generated from another source, which can e.g. be an outdated version of themselves (Tang et al., 2024). However, strictly on-policy RL can be difficult or even impossible to implement in many use cases; it also suffers from sample inefficiency, as each trajectory generated cannot be used for more than a single gradient update. This makes some degree of off-policyness unavoidable, and in fact desirable.
As such, we are interested in the study of off-policy RL techniques, with LLMs finetuning as our motivating example. Various sophisticated methods exist that deal with off-policyness; while classical solutions based on Q-learning or value functions often struggle in the context of language modeling (Zheng et al., 2023), losses based on importance sampling correction have yielded good results, though the variance of the importance ratio can be problematic (Precup et al., 2001; Schulman et al., 2017). In this paper, we consider a simple alternative and study the behavior of a gradient ascent on the expected objective
$J(\pi)={\mathbb{E}}_{y\sim\mu}\left[\log\pi(y)(r(y)-V)\right]$ (1)
as a function of the baseline $V\in\mathbb{R}$, where $\mu$ and $\pi$ are the behavior (sampling) and current policies respectively and $r(y)$ is the reward of trajectory $y$. We call this algorithm Asymmetric REINFORCE (AsymRE). It is asymmetric in the following sense: higher values for $V$ put more emphasis on pushing down the probability of trajectories with low rewards (i.e. “failures”) while mainly ignoring high rewards (i.e., "successes"), whereas a small $V$ results in a more positive approach, where the model mostly increases the probability of successful trajectories while mainly ignoring failures. Our core intuition is that while a model can learn from both its successes and failures while training on-policy, it has less to learn from the failures of another model (i.e., when off-policy), which it might not be likely to produce in the first place, and as such should focus more heavily on the positive examples while training off-policy.
Our key contributions are as follows.
  • In Theorem 4, we show that the AsymRE algorithm applied in a tabular setting converges to a limit policy $\pi^{*}_{\mu,V}$ which we characterize. We show that when the baseline $V$ is smaller than the expected reward $V^{\mu}$ of the behavior policy $\mu$, the limit policy $\pi^{*}_{\mu,V}$ improves upon the behavior policy $\mu$, while maintaining a wide support. When $V$ becomes larger than $V^{\mu}$, a phase transition occurs and the support of $\pi^{*}_{\mu,V}$ shrinks dramatically.
  • In Theorem 4, we study the iterative application of AsymRE in a policy improvement scheme when $V<V^{\mu}$. In combination with Theorem 4, we see that letting $V\geq V^{\mu}$ results in premature convergence to a potentially suboptimal policy.
  • We then verify our findings in a controlled yet rich setting, that of multi-armed bandits.
  • Finally, we validate our results on a larger scale by training Llama 8B and Qwen 3B models on real-world data with AsymRE.
These results confirm our initial intuition that in an off-policy setting, conservatively picking a small baseline is the correct strategy.
2 Related works
Reinforcement learning for LLMs. Reinforcement learning methods are rapidly becoming the dominant paradigm for fine-tuning LLMs on complex tasks, such as mathematical reasoning and coding. It benefits from certain key advantages over supervised training methods: e.g., models can generate their own training samples in the absence of preexisting data of high enough quality. Reinforcement Learning from Human Feedback (RLHF) has emerged as a cornerstone methodology for aligning large language models with human values and preferences (Achiam et al., 2023). Early systems (Ouyang et al., 2022) turn human preference data into reward modeling to optimize model behavior accordingly. DPO (Rafailov et al., 2023) has been proposed as a more efficient approach that directly trains LLMs on preference data. However, as LLMs evolve during training, continuing training on pre-generated preference data becomes suboptimal due to the distribution shift from off-policy data. Thus, the need arises for additional data to be collected during mid-training—a key phase in the iterative fine-tuning of LLMs (Touvron et al., 2023; Bai et al., 2022; Xiong et al., 2024; Guo et al., 2024). One line of works aims to mitigate by merging on-policy and off-policy data to achieve improved performance or alignment (Gu et al., 2017; Feng et al., 2025).
In verifiable domains, recent methods like GRPO (Shao et al., 2024; Guo et al., 2025) have shown strong performance by leveraging binary reward signals. GRPO is a REINFORCE-style algorithm (Williams, 1992) that incorporates negative examples, increasingly recognized as an important ingredient for efficient learning (Ahmadian et al., 2024). At its core, REINFORCE is an on-policy algorithm, with theoretical guarantees only when the reference model matches the trained model, limiting the user’s ability to reuse past data. However, training is rarely truly on-policy: data is generated in a parallel, asynchronous way. Moreover, given the computational cost of rollouts, reusing of trajectories is often desirable.
Off-policy approaches. Algorithms like REINFORCE tend to become unstable with off-policy data (Pang and He, 2021). One way to mitigate these issues is to introduce KL-regularization towards the data-generating policy or early stopping, which effectively diminish the influence of negative trajectory and the amount of off-policy learning. Q-learning (Watkins and Dayan, 1992) is also able to handle off-policy data but is less suitable for deployment in LLMs. Perhaps the most common technique to address off-policy distribution shift is importance sampling in conjunction with clipping in REINFORCE-style algorithms (Schulman et al., 2017; Munos et al., 2016; Espeholt et al., 2018). In practice, though, it is well-known that importance sampling is plagued with excessive variance (Precup et al., 2001). Other approaches to off-policy RL use a consistency condition derived from the explicit solution of the KL-regularized policy optimization problem, which could be enforced on any data, off- or on-policy, see e.g., Rafailov et al. (2023); Richemond et al. (2024); Tang et al. (2025); Cohen et al. (2025). Concurrent work to ours (Roux et al., 2025) introduces an asymmetric variant of importance sampling to speed up learning, noting in passing that the baseline $V$ in REINFORCE plays a role in connection with negative samples in off-policy data.
3 Setting
In reinforcement learning (RL), the task is to maximize the expected reward
$V^{\pi}\stackrel{{\scriptstyle\text{\small{def}}}}{{=}}{\mathbb{E}}_{y\sim\pi}\left[r(y)\right]$ (2)
of a trainable current policy $\pi$ given some reward function $r$. This is typically done by iteratively updating $\pi$ using training samples $(y,r(y))$, where the trajectories $y$ are sampled from some behavior policy $\mu$; here, each trajectory $y$ potentially represents a whole sequence of states (or observations) and actions taken according to the policy, with partial rewards reflected in $r(y)$. If the training samples $y$ are drawn from the current policy, i.e. if $\mu$ is kept equal to the evolving policy $\pi$, it is called on-policy RL; otherwise, it is off-policy RL.
In the absence of a superior behavior policy $\mu$ from which to sample, algorithms that are as on-policy as possible, i.e. in which $\mu$ is kept as close to the trained policy $\pi$ as possible by frequently updating it, is usually preferable for LLM training (Tang et al., 2024). However, on-policy methods suffer from various limitations. Though conceptually simple, they pose delicate engineering problems due to the need for asynchronous cooperation among multiple GPUs, often leading to suboptimal resource utilization. Furthermore, for $\mu$ to remain equal to $\pi$, the algorithm must wait for the current batch of trajectories to be generated and evaluated using the reward function $r$, and for $\pi$ (and $\mu$) to be updated, before a new batch can be produced. In many common settings, such as code generation, agentic interaction with the web or human feedback, rewards can take minutes or even hours to be produced; this is even more true for embedded agents interacting with the physical world. Fully on-policy methods can become extremely inefficient or even infeasible due to this. Finally, a strictly on-policy algorithm can only use the trajectories it generates for a single policy update before having to discard them, whereas one might want to reuse samples generated earlier during training, or even coming from the training of another model. This makes them very sample-inefficient, and consequently compute-inefficient, as trajectories are often costly to generate.
This makes off-policy algorithms an attractive alternative (Rafailov et al., 2023; Richemond et al., 2024; Tang et al., 2025; Cohen et al., 2025); we are in particular interested in off-policy RL with delayed updates, in which the actor policy $\mu$ is simply an outdated version of the current policy $\pi$, and is updated (by setting it equal to $\pi$) every $N$ training iterations of $\pi$. Common off-policy algorithms typically apply importance sampling correction to obtain unbiased estimates of the objective function (Schulman et al., 2017; Espeholt et al., 2018; Tang et al., 2025). Both approaches have drawbacks; in this paper, we investigate a third method for handling off-policy data.
Context-dependent tasks and LLMs.
Context-dependent tasks, such as the ones for which LLMs are designed, add another layer of subtlety to the study of RL algorithms. In the setting above, we considered a single objective function, namely the expected reward from Equation 2. Conversely, the behavior policy $\pi$ of an LLM typically has to generate responses $y\sim\pi(\cdot|x)$ for various prompts $x\sim\rho$ sampled from some task distribution $\rho$, and its goal is to maximize the expected context-dependent reward $\max_{\pi}{\mathbb{E}}_{x\sim\rho,y\sim\pi}\left[r(x,y)\right]$. If each prompt $x$ defined an entirely distinct problem, then whatever RL algorithm we choose to apply would only see a handful of trajectories associated to this problem, and learning would be essentially impossible. This is not so, as the regularities of language and of LLMs’ architectures make it so that similar prompts, such as “Please solve 2 + 3 = ?” and “Compute 2 + 3 = ?”, yield similar problems and conditional distributions, and training on one question helps with the other. This additional structure on the space of conditioning prompts plays a key role in enabling LLM training, yet is rarely discussed in RL literature.
4 Asymmetric REINFORCE
We set ourselves in the general RL setting described earlier in which the goal is to maximize the expected reward $V^{\pi}={\mathbb{E}}_{y\sim\pi}\left[r(y)\right]$ of our current policy $\pi$. For that purpose, one may use a simple on-policy Policy Gradient (PG) algorithm, REINFORCE Williams (1992), which updates the parameters of the policy in the direction
${\mathbb{E}}_{y\sim\pi}\left[\nabla\log\pi(y)\left(r(y)-V\right)\right],$ (3)
where $V\in\mathbb{R}$ is a value baseline and the gradient is taken with respect to some differentiable parameters. In practice, $V$ is often set to be some learnt approximation $V\approx V^{\pi}$ of the value function of the current policy, or to be the empirical average $V=\frac{1}{n}\sum_{i=1}^{n}r(y_{i})$ of the rewards obtained when generating several responses $y_{i}$ (see e.g., Ramesh et al. (2024); Shao et al. (2024)).
In the on-policy setting, the role of the value baseline is simply to offer possible variance reduction of the PG update: the expected behavior of the algorithm during the learning process, as well as its asymptotic performance, are not affected by the choice of this baseline. On the contrary, in the off-policy case, the choice of the baseline impacts the expected behavior of both the algorithm and the asymptotic policy. To analyze this behavior, we consider the simplest policy gradient algorithm, namely REINFORCE, applied to the off-policy setting. We call this algorithm Asymmetric REINFORCE (or AsymRE in short):
Definition 4.1 (Expected AsymRE and AsymRE).
Given a behavior policy $\mu$ and a baseline $V\in\mathbb{R}$, the expected Asymmetric REINFORCE is a gradient ascent in the direction
${\mathbb{E}}_{y\sim\mu}\left[\nabla\log\pi(y)\left(r(y)-V\right)\right],$
The Asymmetric REINFORCE is its stochastic counterpart using samples drawn from $\mu$.
This corresponds to a stochastic gradient ascent on the expected objective:
$J(\pi)={\mathbb{E}}_{y\sim\mu}\left[\log\pi(y)\left(r(y)-V\right)\right].$ (1)
Note that it does not coincide with the expected reward of $\pi$, which is our true objective. As it does not involve any off-policy correction, such as importance sampling, we do not expect this algorithm to converge to an optimal policy in general. However we will see that even off-policy this algorithm offers policy improvement guarantees (under some condition on the choice of the baseline $V$).
Intuition
The baseline $V$ can be understood as a way to control the emphasis put on the good trajectories (those with high rewards) versus the bad ones (low rewards). If $V$ is large, greater (absolute) weight is given to gradient updates from bad trajectories, which are pushed down, and vice versa, if $V$ is low, more weight is given to push-up the good ones. This matters not in an on-policy setting, since the baseline does not introduce bias to the expected dynamics. However, in the off-policy setting, the choice of the baseline in AsymRE changes the expected behavior of the policy gradient and provides us with a way to introduce some asymmetry in our treatment of good versus bad trajectories. In the simplest of settings, in which rewards are binary $r(y)\in\{0,1\}$, letting $V=0$ is equivalent to applying supervised-learning on good trajectories while ignoring the bad ones. Conversely, setting $V=1$ means learning from mistakes only, while ignoring good rewards. Intuitively, we expect a comparatively lower baseline to be more beneficial in an off-policy setting, as the failures of another policy are less informative for the current policy; we explore this intuition in the remainder of this article.
Dynamics of AsymRE in a tabular setting.
Our first result characterizes the dynamics and limits of the AsymRE algorithm in the case of a tabular softmax policy representation.
{restatable}
theoremstablepointsthm[Analysis of expected AsymRE for tabular softmax policies] Let $Y$ be a finite set, $\mu$ be some behavior policy whose support is $Y$, and consider a softmax policy representation $\pi(y)\stackrel{{\scriptstyle\text{\small{def}}}}{{=}}e^{l(y)}/\sum_{y^{\prime}}e^{l(y^{\prime})}$ on $Y$, where the logits $\{l(y)\}_{y\in Y}$ are the policy parameters. We consider the expected AsymRE algorithm with respect to the logits initialized at some $\pi_{0}$ with successive iterates $\pi^{t}_{\mu,V}$ and learning rate $\eta$ as per Definition 4.1. Then the AsymRE algorithm converges to a limit distribution $\pi^{*}_{\mu,V}$. The baseline parameter $V$ controls the nature of the limit distribution:
  • If $V<V^{\mu}$, then $\pi^{*}_{\mu,V}$ is defined by
    $\pi^{*}_{\mu,V}(y)=\frac{(\mu(y)(r(y)-V)-\tau_{\mu,V})^{+}}{V^{\mu}-V},$
    where $\tau_{\mu,V}$ is uniquely characterized by the constraint $\sum_{y\in{\cal Y}}(\mu(y)(r(y)-V)-\tau_{\mu,V})^{+}=V^{\mu}-V,$ and $x^{+}=\max(x,0)$.
  • If $V=V^{\mu}$, then $\pi^{*}_{\mu,V}$ is defined by its support
    $\operatorname{supp}(\pi^{*}_{\mu,V})=\arg\max_{y\in Y}\mu(y)(r(y)-V),$
    and by $\pi^{*}_{\mu,V}(y)/\pi^{*}_{\mu,V}(z)=\pi_{0}(y)/\pi_{0}(z)$ for any $y,z\in\operatorname{supp}(\pi^{*}_{\mu,V})$.
  • If $V>V^{\mu}$, then $\pi^{*}_{\mu,V}$ can charge any of the elements in the set
    $\big{\{}y\,\big{|}\,\min_{z\in Y}\mu(y)(r(y)-V)-\mu(z)(r(y)-V)+V-V^{\mu}>0\big{\}},$
    depending on the initial condition $\pi_{0}$.
As a consequence, $\operatorname{supp}(\pi^{*}_{\mu,V_{1}})\subseteq\operatorname{supp}(\pi^{*}_{\mu,V_{2}})$ when $V_{2}\leq V_{1}\leq V^{\mu}$.
The proof of this theorem can be found in Appendix 7.
Comparison to on-policy REINFORCE:
The result in Theorem 4 is somewhat surprising in the fact that its behavior differs greatly from the on-policy REINFORCE (3). The stable points of the on-policy REINFORCE algorithm are the optimal policies, i.e. those whose support is included in $\arg\max_{y}r(y)$, and as mentioned above, are independent of the choice of the baseline $V$. However, in the off-policy REINFORCE algorithm, the role played by the baseline $V$ is crucial. Not only does it affect training dynamics, but also the final policies. Indeed, the size of the support of the limit policies $\pi^{*}_{\mu,V}$ of AsymRE decreases as $V$ increases, while for $V=\min_{y}r(y)$ the support is $Y$ itself. Thus, the higher $V$, the more deterministic the resulting policies. Moreover, there is an abrupt change of behavior as $V$ becomes equal to or larger than $V^{\mu}$: at this tipping point, the support suddenly shrinks, and generically becomes a singleton. This phase transition has important consequences which we detail below.
Policy improvement scheme with AsymRE:
Let $V<V^{\mu}$, and let us define $\mathcal{T}_{V}$ the operator that takes as input a behavior policy and returns the limit policy of the AsymRE algorithm (as defined in Theorem 4), i.e.:
$(\mathcal{T}_{V}\mu)(y)=\pi^{*}_{\mu,V}(y)\propto(\mu(y)(r(y)-V)-\tau_{\mu,V})^{+}\quad\text{for all}\quad y\in Y.$
Note that this limit policy does not depend on the choice of initial policy $\pi_{0}$. We now consider the repeated application of $\mathcal{T}_{V}$ in a policy improvement scheme, in which the behavior policy $\mu_{t+1}$ of each iteration is the limit policy $\mathcal{T}_{V}\mu_{t}$ of the previous iteration. The expected dynamics of this process are detailed in the following theorem, whose proof can be found in Appendix 7.
{restatable}
theoremdynamicsiteration[Policy improvement dynamics] Let $\mu$ be any policy with support $Y$, and let $V<V^{\mu}$.
  1. 1.
    Each application of the AsymRE algorithm increases the expected reward: $V^{\mathcal{T}_{V}\mu}\geq V^{\mu}$.
  2. 2.
    The sequence of expected rewards $V^{(\mathcal{T}_{V})^{n}\mu}$ converges to some limit expected reward $V^{\infty}$. Let $Y^{\infty}\stackrel{{\scriptstyle\text{\small{def}}}}{{=}}\{y\ :\ r(y)=V^{\infty}\}$. Then the mass of $(\mathcal{T}_{V})^{n}\mu$ concentrates exponentially fast on $Y^{\infty}$, i.e.
    $\sum_{y\not\in Y^{\infty}}((\mathcal{T}_{V})^{n}\mu)(y)\leq c^{n}$
    for some $c<1$.
  3. 3.
    There exists $V_{0,\mu}$ such that $V<V_{0,\mu}$ if and only if the corresponding limit reward is optimal, i.e. $V^{\infty}=\max_{y\in Y}r(y)$.
In practice, the number of training steps in each iteration of a policy improvement scheme is finite; the more steps there are, the more off-policy the training. As the number of steps within an iteration increases, the current policy gets closer to the limit policy, whose support is always smaller than that of the initial policy, and smaller the larger $V$ is. This can lead to the premature elimination of points $y$ with high reward. Hence the more off-policy the training, the smaller the baseline needs to be to prevent premature convergence to a suboptimal solution. Interestingly, it can be shown that one can be over-conservative: letting $V$ be very small results in much slower convergence rates. Thus there is a trade-off between the speed of convergence and the risk of reaching a suboptimal limit policy.
Note that if $V\geq V^{\mu}$, then the first point of the theorem does not hold any more and ${\mathbb{E}}_{y\sim\pi^{*}_{\mu,V}}[r(y)]$ can be smaller than ${\mathbb{E}}_{y\sim\mu}[r(y)]$. Furthermore, the support of the policy shrinks at the first policy improvement iteration (as described in Theorem 4), which makes all subsequent iterations useless.
From the tabular setting to LLMs
As explained at the end of Section 3, LLMs are applied to a variety of conditioning contexts $x$ (the prompts), each of which corresponding to a distinct RL problem. The straightforward adaption of the AsymRE method to this setting is to train an LLM’s policy $\pi$ using the context-averaged version of our loss
${\mathbb{E}}_{x\sim\mathcal{D},y\sim\mu(\cdot|x)}\left[\log\pi(y|x)\left(r(y,x)-V\right)\right],$
where $\mathcal{D}$ is some distribution of prompts. Translating our theoretical findings to the case of LLMs requires some cautiousness. Note first that the expected reward $V^{\mu(\cdot|x)}={\mathbb{E}}_{y\sim\mu(\cdot|x)}\left[r(y,x)\right]$, whose value relative to $V$ conditions the behavior of the AsymRE algorithm, differs for each prompt. A natural solution is to consider instead the context-corrected loss
${\mathbb{E}}_{x\sim\mathcal{D},y\sim\mu(\cdot|x)}\left[\log\pi(y|x)\left(r(y,x)-V^{\mu(\cdot|x)}-\delta V\right)\right],$ (4)
so that the critical value becomes $\delta V=0$ uniformly across the $x$s. Furthermore, though our theorems apply to each of the problems corresponding to the various prompts, the averaged behavior of the model over the distribution of problems is more subtle. At last, our results focus on the training loss of the trained policy, whereas we might be interested in the LLM’s test loss on contexts on which they were not trained.
Nonetheless, we can hope that the regularities of LLMs with respect to prompts can allow us to partially predict the effects of training on certain prompts on the LLM’s policy conditioned by other prompts. In particular, our results suggest that letting $V$ be equal to or larger than 00 in the context-corrected loss in Equation 4 can lead to dramatic decrease in the diversity of answers, corresponding to the collapse of the policy’s support predicted by our theorems. While this may not necessarily penalize the expected reward of a tabular policy, it is detrimental to a language model; in particular, it poses a risk of severe overfitting of the training set, and consequently of poor test results.
5 Experiments
We validate our findings first in a controlled setting, then with LLMs on real-world data.
5.1 Bandits
We illustrate our findings in a simple bandit setting. There are 100 arms. The expected reward $r(y)$ of each arm $y$ is chosen uniformly randomly in $[0,1]$. We sample from a non-uniform behavior policy $\mu$ defined as a softmax of the logits $l(y)=y/10$, where $y$ is indexed from 00 to $99$. The current policy $\pi_{t}$ is a softmax of its logits $l_{t}(y)$, initialized as $\pi_{0}=\mu$, and its logits are updated according to the expected AsymRE algorithm with learning rate $1$:
$l_{t+1}=l_{t}+\nabla_{l}{\mathbb{E}}_{y\sim\mu}\left[\nabla\log\pi_{t}(y)\left(r(y)-V\right)\right].$
(a) Expected reward ${\mathbb{E}}_{y\sim\pi_{t}}[r(y)]$ of the AsymRE algorithm as a function of the training iteration $t$ and baseline $V$.
(b) The support of the final policy $\pi^{*}_{\mu,V}$ is shown by the yellow dots; each dot corresponds to an arm (sorted by increasing reward values). The support suddenly drops to a single atom when $V$ crosses $V^{\mu}\approx 0.54$.
Figure 1: Expected rewards and supports of the policies in the bandits experiments
Expected reward and support with AsymRE
In Figures 1(a) and 1(b), we report the expected reward ${\mathbb{E}}_{y\sim\pi_{t}}[r(y)]$ of the expected AsymRE algorithm. We observe that the performance of the algorithm tends to improve as the value of the baseline $V$ gets closer to some $V_{0,\mu}<V^{\mu}\approx 0.5405$. However, the performance of all policies is upper bounded by a sub-optimal value $\approx 0.89$, which is lower than the optimal expected reward ($\max_{y}r(y)=0.999$ for this run).
Although these experiments may seem to support the choice of selecting a high value of the baseline, Figure 1(b), where we plot the support of the final policy $\pi^{*}_{\mu,V}$ as a function of $V$, shows a complete loss of diversity in the resulting policies for high values of $V$, making future improvements in a policy iteration scheme impossible. This is also supported by the drastic decrease in entropy for high values of $V$ reported in Figure 6, in the Appendix. We observe a phase transition for $V$ around $V^{\mu}$. For $V<V^{\mu}$, the size of the support decreases as $V$ increases, as predicted in the Theory Section 4. The policy stays supported by a large number of atoms even when $V\approx V^{\mu}$, as long as $V<V^{\mu}$. But as soon as $V\geq V^{\mu}$, the support of the optimal policy $\pi^{*}_{\mu,V}$ becomes a singleton: $\arg\max_{y}\mu(y)(r(y)-V)$.
Thus, when $V\geq V^{\mu}$, there is a risk that our off-policy REINFORCE algorithm might produce policies that lose their diversity. In addition, notice that when $V$ is slightly larger than $V^{\mu}$, the reward of the corresponding limit policy is not as large as the reward of the best arm in the support of the limit policy for $V$ slightly smaller than $V^{\mu}$. Thus the potential of finding this superior arm is lost forever when $V$ crosses $V^{\mu}$.
Policy improvement with AsymRE
We study a policy improvement scheme over $40$ iterations, where at each iteration we apply the AsymRE algorithm for $500$ steps using the policy obtained at the end of the previous iteration as both our behavior policy and starting current policy. In Figure 2, we show the expected reward of the corresponding policies. We notice that when $V<V^{\mu}$, for every policy iteration step, the performance of the improved policies $\mathcal{T}_{V}\mu_{t}$ are better than the corresponding behavior policies $\mu_{t}$, as predicted by Theorem 4. However, for $V\geq V^{\mu}$, we have already seen in Figure 1(a) that the policy converges to a deterministic and suboptimal policy. Here we see that iterating this algorithm does not permit to recover from this premature suboptimal convergence, regardless of the value of $V\geq V^{\mu}$. We also observe that while they do not plateau as dramatically, trajectories corresponding to $V$ only slightly smaller than $V^{\mu}$ (such as $0.525$ and $0.54$) converge to a less favorable limit policy than for $V\leq 0.5$.
Figure 2: Performance of the current policies for 40 iterations of policy improvement with AsymRE. Each iteration is 500 steps. The curve corresponding to $V\approx V^{\mu}$ is in red.
5.2 Large Language Models
We validate our findings by training LLMs on reasoning tasks.
Experimental setup
We train Llama-3.1-8B-Instruct (Dubey et al., 2024) (which we refer to as Llama 8B) and Qwen2.5-3B-Instruct (Yang et al., 2024) (which we refer to as Qwen 3B) with the AsymRE objective
$J(\pi)=\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\mu(.|x)}\Big{[}\frac{1}{G}\sum_{i=1}^{G}(r(y_{i},x)-(\hat{V}+\delta V))\log(\pi(y_{i}|x))\Big{]},$
where $x$s are prompts from the dataset $\mathcal{D}$, each $y_{i}$ is a full sequence of tokens auto-regressively sampled from the LLM $\mu$, and $\hat{V}=\frac{1}{G}\sum_{i=1}^{G}r(y_{i},x)$ is the empirical estimate of the average reward $V^{\mu(\cdot|x)}$. This loss is the empirical approximation of the context-corrected loss in Equation (4). We place ourselves in the delayed updates setting, i.e. the behavior policy $\mu$ is simply an outdated version of the current policy $\pi$, which we update every $N$ training steps; this corresponds to a policy improvement scheme where each iteration lasts $N$ steps. We set the number $G$ of samples per prompt to $8$. We train and test on the MATH dataset (Hendrycks et al. (2021), $12.5$k high school level problems). The reward of a trajectory is $1$ if the answer is correct, and $-1$ otherwise. Additional details regarding the experimental setup are given in Appendix 8.
Impact of the baseline $\delta V$
Figure 3: Training dynamics of Llama 8B on the MATH dataset (results are averaged over $3$ seeds, and a moving average with a window of size $3$ is applied). The behavior policy is updated every $N=250$ training steps.
The train and test losses for various baselines $\hat{V}+\delta V$ are shown in Figure 3. Similarly as what we observe with the bandits experiment (Figure 2), while the baseline $\delta V$ is smaller than the critical value 00, the model’s training accuracy tends to increase with $\delta V$. This is reflected in the test accuracy as well, though less strongly. When the baseline $\delta V$ reaches, then passes 00, a phase transition occurs and both the training loss and the test loss suffer a catastrophic collapse. The larger $\delta V\geq 0$, the faster this collapse occurs (for $\delta V=0$, the collapse occurs at the very end of our training run). This corresponds in the theory (and in the bandit setting) to the regime where the limit policy becomes deterministic: while this can result in a good expected reward in the single-problem bandit setting, it leads to a drastic decrease in accuracy in the LLM setting, where the model must retain diversity to answer the range of problems with which it is tasked. This loss of diversity can be observed in the entropy of the models, in Figure 7 in the Appendix.
Figure 4 confirms our claim that letting $\delta V$ be slightly smaller than 00 results in greater training stability: we compare $7$ independent training runs with $\delta V=0$ and $7$ training runs with $\delta V=-0.1$, and we find that all runs with $\delta V=0$ collapse.
Our additional results on Qwen 3B (Figure 8) in Appendix 9 support the same conclusions.
Figure 4: Training dynamics of Llama 8B on the MATH dataset for two values of the baseline, $\delta V=0$ and $\delta V=-0.1$, and $7$ independent runs for each value. The behavior policy is updated every $N=250$ steps. We observe a systematic collapse when $\delta V=0$.
5.3 Discussion
These experiments corroborate our theoretical findings and can be summarized as follows:
  • Letting the baseline $V$ be strictly smaller than the expected reward $V^{\mu}$ of the behavior policy $\mu$ makes the training stable, even when off-policy, and offers a monotonic improvement of the policies.
  • Performance tends to improve as $V$ grows closer to $V^{\mu}$ while remaining strictly smaller.
  • As soon as $V$ passes $V^{\mu}$, the policies become more deterministic. In the context of bandits, this leads to a suboptimal convergence of the policy improvement scheme. In the context of LLMs, this leads to a collapse of the training and test loss.
While further large scale experiments are needed to draw definitive conclusions regarding the training of LLMs, our findings suggest that adding a small conservative correction $\delta V\approx-0.1$ in the advantage term of the training objective
$\mathbb{E}_{x\sim\mathcal{D},\{y_{i}\}_{i=1}^{G}\sim\mu(.|x)}\Big{[}\frac{1}{G}\sum_{i=1}^{G}(r(y_{i})-(\hat{V}+\delta V))\log(\pi(y_{i}|x))\Big{]}$
might result in greater training stability, and help prevent the collapses often observed in RL training.
6 Conclusion
We have thoroughly studied the behavior of our simple off-policy algorithm AsymRE with respect to the baseline $V$, both theoretically and experimentally. We have confirmed our claims that in an off-policy setting, selecting a baseline slightly smaller than the expected reward of the behavior policy results in superior asymptotic performances, and can help alleviate the risk of training collapse. Such a choice of baseline corresponds to putting more emphasis on positive training examples, and less on negative ones, which is intuitively reasonable.
Limitations and possible improvements
While we have analyzed in depth the role of the baseline $V$ in our simple AsymRE method, it would be interesting to extend our study to more sophisticated losses, which might e.g. include importance ratio correction and KL-divergence regularization. Finally, the possibility to leverage off-policyness by training over the same samples for multiple epochs was one of our initial motivations: quantifying the potential run-time and compute gains and the impact on model performance of such a scheme is an exciting future research direction.